OpenStudy (anonymous):

volume of a cylinder with a ratio of 4/2 or 2/1 the volume of the smaller cylinder is 70 my answer either 560 or maybe 280 sing another formula im leaning more towards 560 ...

OpenStudy (anonymous):

using*

OpenStudy (lgbasallote):

4:2 is the ratio.so obviously the 2 is the smaller volume so your proportion will become... 4:2 = x: 70 right? so... \(\Huge \frac{4}{2} = \frac{x}{70}\) what do you do now?

OpenStudy (lgbasallote):

the ratio 4:2 is the ratio of the volume of the larger cylinder to the volume of the smaller cylinder right?

OpenStudy (anonymous):

well /2 can be 2/1 then should be 2 x 2 x2 =8/1 8x70=560

OpenStudy (anonymous):

4/2

OpenStudy (lgbasallote):

wait..where'd you get 2 x 2 x 2 again?

OpenStudy (anonymous):

its either 8/1 or 4/1

OpenStudy (anonymous):

from 4/2 it can be 2/1 right then you have o multiply its a rule the 2 x 2 x2 =8/1

OpenStudy (anonymous):

however im unsure if its times 3 or times 2 sense its 4/2

OpenStudy (lgbasallote):

8/1 is NOT equal to 4/1 what they meant was that 4/2 can be simplified to 2/1

OpenStudy (anonymous):

yeah but then you multiply the top number of the alrdy simplified number

OpenStudy (lgbasallote):

you cannot multiply them...you're not being asked to multiply..they just simplified it for you

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

so that's how you'll get your x... \(\Huge \frac{4}{2} = \frac{x}{70}\) or \(\Huge \frac{2}{1} = \frac{x}{70}\) just cross multiply